Corrections to “The Proof of a conjecture of Simion for certain partitions”, M. Hildebrand, Disc. Math. 224 (2000), 139-150

On page 141, in the proof of Proposition 2, there is a denominator which reads

1 − \binom{j}{n + 1}

but should read

1 − \frac{j}{n + 1}

instead.

The author regrets this error.